Circular No. 5856

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 FW

Further precise positions have been reported as follows:

1993 UT α_{2000} δ_{2000} R Observer
May 17.35082 12^{h}23^{m}58.^{s}68 −2^{o}38′22′′.6  Hainaut
17.39862 12 23 58.55 −2 38 22.0  Byun
18.01676 12 23 56.96 −2 38 09.8  _{”}
June 25.05761 12 23 19.44 −2 32 50.2 23.8 Marconi

Y.-I. Byun (Mauna Kea). 2.2-m reflector + CCD. Measurers J. Luu and

D. Jewitt. O. Hainaut and G. Marconi (European Southern Observatory). 1.5-m Dan-

ish reflector and 3.6-m NTT + CCD. Measurers R. M. West and O. Hainaut.

The orbit of this object is still very indeterminate, but the following result, derived by the undersigned from the 11 available observations (cf. IAUC 5730, 5796), suggests the low orbital eccentricity and general similarity to the orbit of 1992 QB_{1} (cf. IAUC 5855). The object is assumed to have been at perihelion at discovery. Further observations are very necessary, however, and observers should attempt a recovery as soon as the object returns to the morning sky and is still in direct motion.

T = 1993 Mar. 28.180 TT  ω = 359.^{o}501
e = 0.04066  Ω = 187.914 }2000.0
q = 42.1448 AU  i = 7.745
a = 43.9311 AU n^{o} = 0.0033849 P = 291.2 years
1993/94 α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Nov. 19 12^{h}34.^{m}13 − 3^{o}37.^{′}7 42.806 42.145 47.^{o}5 1.^{o}0 23.4
29 12 34.79 − 3 41.6 42.668 42.145  57.4 1.1 23.4
Dec. 9 12 35.32 − 3 44.7 42.513 42.145  67.4 1.2 23.4
19 12 35.72 − 3 46.9 42.347 42.145  77.5 1.3 23.4
29 12 35.98 − 3 48.1 42.175 42.145  87.6 1.3 23.4
Jan. 8 12 36.08 − 3 48.4 42.001 42.145  97.8 1.3 23.4
18 12 36.03 − 3 47.8 41.831 42.145 108.0 1.3 23.4
28 12 35.83 − 3 46.2 41.671 42.145 118.2 1.2 23.4
Feb. 7 12 35.49 − 3 43.7 41.525 42.145 128.4 1.1 23.4
17 12 35.03 − 3 40.4 41.398 42.145 138.7 0.9 23.4
27 12 34.47 − 3 36.4 41.294 42.145 148.9 0.7 23.3

1993 September 4 Brian G. Marsden